OpenStudy (anonymous):

g(x) = x + 1 - e^x find the maximum value of g(x0, and hence show that x + 1 (lesser than or equal to) e^x

OpenStudy (anonymous):

aw hi parth :)

Parth (parthkohli):

Hey. I think I know how to answer this :-)

OpenStudy (anonymous):

ok SHOOT NOW. so i got the max point (e^x = 1)

OpenStudy (anonymous):

but what do i do after that?

Parth (parthkohli):

Since \(e^x\) is much more rapidly increasing than \(x + 1\) is, we can already guess that the max value of \(g(x) \) is at the point \(x = 0\)

Parth (parthkohli):

\[g(0) = 0 \le e^0 = 1\]

OpenStudy (anonymous):

but i got e^x = 1. what do i do with this??

Parth (parthkohli):

that is the max point.

OpenStudy (anonymous):

ok. now what?

Parth (parthkohli):

Since the maximum point, \(g(0)\), is positive, we can say that \(x + 1 \le e^x\)

Parth (parthkohli):

for all \(x\)

OpenStudy (anonymous):

sorry, i don't fully understand. Can you explain it a little more??

OpenStudy (anonymous):

just the last part: how we got to the conclusion.

Parth (parthkohli):

No, I meant to say something else. I mean that the max value is non-negative.

Parth (parthkohli):

Let us assume \(f(x) = x + 1\) and \(g(x) = e^x\)

OpenStudy (anonymous):

ok, please explain. I'm going for dinner, i'll be back in 30 mins ish, so don't worry if i don't respond. anyhoo, thank you SOO much in advance :P

Parth (parthkohli):

Or rather \(f(x) = x + 1\) and \(p(x) = e^x\)

OpenStudy (anonymous):

i REALLy need help :P kay bye.

Parth (parthkohli):

Suppose that we have a number \(x_1\) where \(f(x) - g(x) \) is the maximum, and \(f(x_1) - g(x_1) > 0\) and both \(f(x),g(x)\) are increasing. Then we can say that \(g(x) \ge f(x) \) for all \(x\)

Parth (parthkohli):

Can you see how this works?

Parth (parthkohli):

OK, see this: If \(g(x)\) increases as \(x\) increases, then \(x + 1\) is increasing faster than \(e^x\) and so \(x + 1 \ge e^x\). If \(g(x)\) decreases as \(x\) increases, then \(e^x\) is increasing faster than \(x + 1\) and so \(e^x \ge x + 1\). If \(g(x)\) stays the same all the time, then \(e^x = x + 1\). Do you see where I am getting?